aluminum maker Alcoa has a beta of about 1.86​, whereas Hormel Foods has a beta of

0.38. If the expected excess return of the market portfolio is 3​%, which of these firms has a higher equity cost of​ capital, and how much higher is​ it?

  1. Alcoa's equity cost of capital is ___%?
  2. Hormel's equity cost of capital is ____%?
  3. Therefore, ____?____ has the higher equity cost of capital by __?___ percentage points.
Answer #1

Equity cost of capital=Rf+Beta(Rm-Rf)

where Rf=risk free rate

Rm=market rate of return

Rm-Rf=expected excess return of the market portfolio(market risk premium)

Thus for Alcoa cost=Rf+(1.86*3%)=Rf+5.58%

For Hormel Foods,cost=Rf+(0.38*3%)=Rf+1.14%

Hence higher cost of capital is of Alcoa by (5.58-1.14)=4.44% as the Risk free rate remains the same.